# Out of spuriousity: Improving robustness to spurious correlations without group annotations

### v1

@@ Abstract
Machine learning models are known to learn spurious correlations, i.e., features having strong relations with class labels but no causal relation. Relying on those correlations leads to poor performance in the data groups without these correlations and poor generalization ability. To improve the robustness of machine learning models to spurious correlations, we propose an approach to extract a subnetwork from a fully trained network that does not rely on spurious correlations. The subnetwork is found by the assumption that data points with the same spurious attribute will be close to each other in the representation space when training with ERM, then we employ supervised contrastive loss in a novel way to force models to unlearn the spurious connections. The increase in the worst-group performance of our approach contributes to strengthening the hypothesis that there exists a subnetwork in a fully trained dense network that is responsible for using only invariant features in classification tasks, therefore erasing the influence of spurious features even in the setup of multi spurious attributes and no prior knowledge of attributes labels.
@@ 1 Introduction
Deep neural networks tend to learn spurious attributes, especially when spurious attributes have a positive relation with target labels and are easier to learn which is known as simplicity bias (Shah et al., 2020). This phenomenon leads to the scenario that deep neural networks under-learn invariant information and then lose generalization ability and perform poorly in the data group that does not hold spurious correlations. Spurious features may occur naturally and some of them, such as image background, can be easily detected, whereas other implicit spurious features, such as texture (Geirhos et al., 2018), superficial statistics (Wang et al., 2019) or frequency bias defined by Fourier domain (Wang et al., 2022; 2023) requires more effort. An example of a natural spurious correlation of background is the cows and camels classification task, where the background desert is spuriously correlated with the class camels and grass positively correlates with cows (Beery et al., 2018). These features seem to be easy to learn for deep networks, but relying on them makes networks perform poorly in minority groups such as camels on grass background.
@@
To directly improve the performance of minority groups, mitigation strategies use group label information during training to minimize the worst-group loss (Sagawa et al., 2020a) or re-direct the classification distribution by retraining part of deep networks (Kirichenko et al., 2023). However, group labels can only be collected if one has prior knowledge of the spurious correlations within the dataset, and even in that case, the label annotation process requires much time and effort and might be impossible in a realistic scenario. For example, Wang et al. (2022) defined a frequency shortcut type based on Fourier domain. However, the detection of this type of shortcut is heavily influenced by the order of frequency removal, and their observations are limited to texture-based shortcuts (Wang et al., 2023). To overcome the limitation of collecting group labels, a scope of work leverages hard cases in the training data and defines a weighted loss for retraining models (Liu et al., 2015; Nam et al., 2020; Zhang et al., 2022; Park et al., 2023). Hard cases are those instances that are misclassified by a model trained with empirical risk minimization (ERM). This definition is based on the assumption that models that are influenced by spurious correlations fail to classify instances where the spurious correlation is not present. However, this assumption is not guaranteed to hold. Further, hard cases are not available in the case of zero training loss, i.e., when the model fully fits the training data. Then, careful early-stopping control is required to obtain a sufficient amount of hard cases without fully detoriating the ERM training.
@@
To overcome the limitation of the assumption that ERM misclassified instances are hard and spurious-free cases, we propose Pruning Spurious Correlations (PruSC) that defines potential spurious groups based on the representation of instances in latent space. Our main motivation is that if the ERM models have spurious correlations, we assume that instances with the same spurious features lie close together in the representation space. That means we can identify groups of intances that have the same spurious feature by applying unsupervised clustering on the output of the penultimate layer. Consequently, to reduce the effect of spurious correlations, we distort the clusters of spurious features in feature space. To do that, we introduce a novel application of supervised contrastive loss for training a subnetwork by defining positive and negative anchors based on both target labels and their position in the feature space. By the distortion of spurious feature clusters in representation space, the subnetwork cannot rely on learning and grouping samples based on similar spurious attributes and hence is less influenced by spurious correlations.
PruSC does not require prior knowledge about spurious correlations and is even applicable when more than one spurious attribute is present or the attributes are not trivial for human to annotate (e.g., texture or fluency bias (Geirhos et al., 2018; Wang et al., 2023)). In addition, clusters of spurious attributes form in the representation space as soon as models rely on spurious correlations, i.e., also when models have zero training loss. Hence, our approach to mitigate spurious correlations can be directly applied to a fully trained ERM model
In summary, our contributions are:
  • We propose a subnetwork pruning approach PruSC to mitigate the influence of spurious correlations in deep neural networks that does not require group annotations or prior knowledge of spurious features, not even in the validation set for fine-tuning hyper-parameters.
  • We propose an effective way to construct contrastive learning batches based on representation clustering.
  • Our approach can be applied directly to a trained network without any control requirements (e.g., misclassified cases defined from early stopping).
  • Our approach is robust to multiple spurious correlations.
@@
We evaluate our approach on the CelebA dataset (Liu et al., 2015) with potential multiple spurious attributes and the Skin Cancer ISIC dataset (Codella et al., 2019), a realistic image collection used to detect melanoma and an artificial benchmark WaterBird (Sagawa et al., 2020a). Our method outperforms all approaches that do not require group annotations, even for tuning hyper-parameters, and is comparable with methods that require group labels during training. Our code is available at: https://github.com/pquynhle/spurious-free-subnetwork
@@ 2 Related Work
Approaches to mitigate spurious correlations can be broadly classified in two main categories: having prior knowledge of spurious correlations (such as group labels or human annotations) or not having that information.
With full information about spurious correlations on all instances, data-centric approaches like UV-DRO (Srivastava et al., 2020) and LISA (Yao et al., 2022) augment the training data with additional samples to eliminate the spurious correlation. Group-DRO (Sagawa et al., 2020a) instead utilizes the existing group labels directly to minimize the loss for worst-group performance.
@@
Requiring only a small subset of group labeled training data, Barack (Sohoni et al., 2022) predicts other instances’ groups and DFR (Kirichenko et al., 2023; Izmailov et al., 2022) retrains part of the model with a small, spurious-free dataset. Prior knowledge of spurious correlations has the advantage that training regimes can directly mitigate undesired attribute influence or relations in the data and consequently yields high accuracy, both on average and in the minority group. However, those group annotations or human-in-the-loop annotations are costly and hard to attain in practice. In particular complex spurious attributes, such as as fluency bias (Wang et al., 2023) or texture and shape bias (Geirhos et al., 2018) remain challenging, as they are hard to recognize for humans.
@@
When group annotations are not available, (Sohoni et al., 2020; Seo et al., 2022) and Creager et al. (2021) focus on automatically inferring appropriate groups before training a robust model (e.g., Group-DRO) with pseudo-group labels. Multiple methods follow a two stage training approach, where the first stage commonly serves to identify hard cases from ERM training. These hard cases are instances that the ERM model fails to classify, presumably because ERM relies on spurious correlations but in those instances the correlation is not present (also called bias-conflict). Approaches differ in the second stage: (Liu et al., 2021; Nam et al., 2020) upweight hard cases in an additional training run, while (Yaghoobzadeh et al., 2021) exclusively fine-tune hard cases. CnC (Zhang et al., 2022) and DCWP (Park et al., 2023) utilize the set of hard cases for contrastive learning to define contrastive batches and retrain the whole network (CnC) or to extract a subnetwork (DCWP) with contrastive loss. Different from two-stage methods, DeDiER (Tiwari et al., 2024) uses knowledge distillation to find over-confident predictions in early network layers $d$ then leverages this signal into a graded adjustment of distillation loss on a per-instance level. SELF (LaBonte et al., 2024) finds a small set of hard cases by measuring the degree of disagreement in predictions between an ERM and an early-stopped ERM model, then applies last-layer fine-tuning with this set.
The effectiveness of methods that rely on ERM misclassifications for hard cases depends on both quantity and quality of misclassifications. In cases when ERM models almost perfectly fit the training data (training accuracy close to $100\%$), there are not enough hard cases (i.e., misclassifications) and these methods fail. To obtain a sufficiently large set of hard cases, careful control of early-stopping regularization or choice of layer $d$ (DeDiER) is required. Hence, those methods can not directly mitigate spurious correlations of a trained ERM model unless trainers have access to the whole training process of that model.
Modular Networks.
Modular networks or subnetworks have previously been investigated regarding statistics and connection clusters (Filan et al., 2020). Csordás et al. (2020) proposes a technical training approach for subnetworks in terms of functional modular, a subnetwork highly correlated to a designed task. For example, Csordás et al. (2020) investigates whether there exists a subnetwork responsible purely for one class. Using the same technique, Zhang et al. (2021) utilizes the functional modular to solve out-of-distribution generalization by probing subnetworks serving different functional subtasks from a pre-trained ERM.
@@ 3.1 Problem Setting
Assume a dataset $\mathcal{D}$ consisting of data samples belonging to groups $g\in\mathcal{G}$ and that the groups have different levels of biases correlated to an input feature of $x$ and the label of $y$, i.e., a spurious correlation (Sagawa et al., 2020b). We set each group $g\in\mathcal{G}$ to be defined by the combination of the label $\mathcal{Y}$ and a corresponding bias attribute $\mathcal{A}$. Hence, $g$ follows $\mathcal{G}=\mathcal{A}\times\mathcal{Y}$, and we say that there exists a spurious correlation between $\mathcal{A}$ and $\mathcal{Y}$.
A model trained with ERM on the class labels might be able to separate the classes well but rely on spurious features to do so, resulting in a mismatch between the intended and model-learned solution (Geirhos et al., 2020). Spurious features typically appear in the majority of the dataset. Thus, by using these features, an ERM model can minimize the training loss for the expected average population but still show high errors over the minority group, which lacks these spurious features. Furthermore, in cases where the spurious features are simple to learn (e.g., background instead of the object), ERM prioritizes learning those features over more complex ones and then loses generalization ability. Our goal is to mitigate spurious correlation learning by addressing both, the statistical (highly imbalanced group distributions) and the geometric issues (tendency to learn simpler features).
@@ 3.2 Motivating Example
To illustrate the problem and motivate our approach, we revisit a simple 2-D classification task with spurious features (Pezeshki et al., 2021). The two moons dataset $\mathcal{D}$ consists of input pairs $(x,y)$, where $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ and $y\in\{0,1\}$ (cf. Fig. 1). The dataset contains the artificial direct relation between $x_{1}$ and the corresponding label $y$.
@@
We train three different models on this dataset, showing the effect of i) smaller networks and ii) our contrastive loss formulation (Sec. 5.2.1). A simple feedforward network (5 ReLu layers, 500 hidden units per layer, 100 epochs, cross-entropy loss) can separate the two classes very well (Fig. 1, left). The decision boundary is nearly linear because the cross-entropy loss optimizes for class separation and the model can rely on the correlation between the x-axis values of $x$ and $y$. We train the same network architecture with random masks (50% of weights in each layer) and use the same fraction of random weights for testing, i.e., we randomly prune the network to 50% of its weights. We observe curved decision boundaries (Fig. 1, center), suggesting that a smaller network within a dense network is less influenced by spurious correlations. Combining masked training and our contrastive loss (Sec. 5.2.1, we obtain a curved decision boundary with a large margin (Fig. 1, right), suggesting that the model uses information from both, the x and y coordinate.
@@
Figure 1: Simple network on the two moons dataset. Decision boundaries (black lines) depend on model capacity and loss. Left: standard ERM with cross-entropy loss with a nearly linear decision boundary and a strong dependence on spurious feature (x-coordinate). Center: Pruned network (masking) using only 50% of the weights at test time shows less dependency on the spurious feature. Right: Masking, and using our contrastive loss (cf. Sec. 5.2.1) results in non-linear decision boundaries with large margins.
Implications.
The above example strengthens two hypotheses for mitigating spurious correlations.
@@
First, capacity seems to have an influence. Both, the dense network and the smaller subnetwork with randomly pruned connections perform well on the dataset, but the pruned network has better decision boundaries, i.e., depends less on the spurious feature. However, that does not suggest to train a small architecture from scratch as larger models have a higher ability to learn information and improve performance beyond the bias-variance regime (Nakkiran et al., 2021). Instead, it suggests to first train a large network with large enough capacity to encode all relevant information and then, post-training, extract a subnetwork from it that does not rely on spurious correlation.
The post-training approach also matches our realistic scenario that the existence of spurious correlations becomes evident only after training (and possibly later in the lifetime of a model). Thus, the extraction of a subnetwork without spurious correlations is less expensive than retraining the whole model from scratch. This observation serves as our main motivation.
Second, our contrastive learning scheme improves decision boundaries by learning more information from training data than just the simplicity bias features. Our batches of contrastive instances serve as guidance for models to know which feature relations should be less focused on.

### v2

@@ Abstract
Machine learning models are known to learn spurious correlations, i.e., features that have strong correlations with class labels but no causal relationship. Relying on these correlations leads to poor performance in data groups that do not contain these correlations, and poor generalization. Approaches to mitigate spurious correlations either rely on the availability of group annotations or require access to different model checkpoints to approximate these group annotations. We propose PruSC, a method for extracting a spurious-free subnetwork from a dense network. PruSC does not require prior knowledge of the spurious correlations and is able to mitigate the effect of multiple spurious attributes. Specifically, we observe that Empirical Risk Minimization (ERM) training leads to clusters in representation space that are induced by spurious correlations. We then define a supervised contrastive loss to extract a subnetwork that distorts such clusters, forcing the model to learn only class-specific clusters, rather than attribute-class specific clusters. Our method outperforms all annotation-free methods, achieves worst-group accuracy competitive with methods that require annotations and can mitigate the effect of multiple spurious correlations. Our results show that in a fully trained dense network, there exists a subnetwork that uses only invariant features in classification tasks, thereby eliminating the influence of spurious features.
@@ 1 Introduction
Deep neural networks tend to learn spurious correlations or shortcuts, i.e., misleading heuristics that work for the majority of training samples, but do not always hold (Sagawa et al., 2020a). Spurious attributes or features are features that have no causal relationship with class labels. The existence of spurious features in a data set itself is not harmful, as long as a model does not rely on them for the classification. However, spurious features are often easy to learn, and if there is a strong correlation with class labels, i.e., a spurious correlation, models tend to rely on these spurious correlations as shortcuts, because learning these correlations is sufficient to minimize the overall loss (simplicity bias) (Shah et al., 2020). This means, that in the presence of spurious correlations, deep neural networks under-learn invariant information and then lose their ability to generalize and perform poorly on data that does not contain the spurious correlations. [^1] Shortcut learning can also occur on (subsets of) non-spurious features that do not fully describe the relationship between features and class labels. Ideally, we aim for robust models that rely on invariant features, i.e., they do not rely on a particular distribution of the training data. Spurious features occur naturally, and some of them, such as image background, can be easily detected, while others, such as texture (Geirhos et al., 2018), superficial statistics (Wang et al., 2019), or frequency bias (Wang et al., 2022; 2023), require more effort. An example of a natural spurious correlation of the background is the cows and camels classification task, where the background desert is spuriously correlated with the class camels and grass is positively correlated with cows (Beery et al., 2018). These features seem to be easy to learn for deep networks, but relying on them makes networks perform poorly in minority groups such as camels on grass.
@@
To improve the performance for minority groups, mitigation strategies use group label information, obtained either by manual labeling or by approximation using misclassification information. Ground-truth group labels have been used during training to minimize the worst-group loss (Sagawa et al., 2020a) or to retrain only the last layer (Kirichenko et al., 2023). However, obtaining group labels either requires prior knowledge of the spurious attributes and a time-consuming annotation effort, or may be impossible for certain types of attributes. For example, neural networks tend to rely on low frequency components for regression tasks (Rahaman et al., 2019). Wang et al. (2022; 2023) showed that classification models also rely on frequency (both texture-based and shape-based). However, frequency attributes are only implicit and not visible to the human eye, unlike explicit spurious attributes, such as background (e.g., grass or desert). A common proxy estimate for missing group labels is to use misclassified examples from the training data (i.e., hard cases), based on the assumption that models that are influenced by spurious correlations fail to classify instances where the spurious correlation is absent (Liu et al., 2015; Nam et al., 2020; Zhang et al., 2022; Park et al., 2023). However, this assumption is not guaranteed to hold. Furthermore, in the case of zero training loss, where the model fits the training data perfectly, hard cases are not available. Careful control of early stopping is then required to obtain a sufficient number of hard cases without degrading the Empirical Risk Minimization (ERM) training. Consequently, these methods also require access to full training information and are not applicable to fully trained models.
@@
In this paper, we propose Pruning Spurious Correlations (PruSC), a method for extracting a spurious-free subnetwork from a fully trained dense network that requires no prior knowledge of the spurious correlations, i.e., no ground-truth group annotations. PruSC defines potential spurious groups based on the representation of instances in latent space. Based on the observation that instances with the same spurious attributes lie close together in representation space, we can identify groups of instances with the same spurious feature. We apply clustering to the representations of the penultimate layer and introduce a novel application of supervised contrastive loss to train a subnetwork such that clusters (likely) induced by spurious attributes are distorted. By distorting the clusters induced by the spurious feature, the subnetwork cannot rely on learning and grouping samples based on similar spurious attributes and is, therefore less influenced by spurious correlations. In summary, the contributions of this paper are:
  • We present PruSC, an approach to mitigate the effects of spurious correlations. Our method relaxes the assumptions of previous work: PruSC does not require prior knowledge of spurious correlations, and it does not require any group annotations in either the training or the validation set.
  • We show how to extract a subnetwork optimized for unlearning data manifolds induced by spurious attributes. We base our solution on contrastive learning, and introduce an effective way to sample contrastive batches based on representation clustering.
  • We evaluate PruSC on benchmark datasets and show that it has superior worst-group accuracy to annotation-free approaches and is competitive with approaches that require annotations.
  • We show that PruSC is robust to multiple spurious correlations and achieves the largest improvement over all spurious attributes.
@@
We evaluate our approach on the CelebA dataset (Liu et al., 2015) with potentially multiple spurious attributes, the ISIC skin cancer dataset (Codella et al., 2019), a realistic collection of images used for melanoma detection, and the artificial benchmark Waterbirds (Sagawa et al., 2020a). Our method outperforms all approaches that do not require group annotations across all benchmarks and has the highest overall score on 2 out of 3 benchmarks. Our code is available at: https://github.com/aix-group/prusc.
@@ 2 Related Work
In this section, we review related work on mitigating spurious correlation. Approaches to mitigate spurious correlations can be broadly categorized based on assumptions: methods that require prior knowledge of spurious correlations, such as group labels or human annotations, and methods that do not require prior knowledge.
@@
With prior knowledge of spurious correlations, training regimes can directly mitigate undesired attribute influences or relations in the data and consequently yield high accuracy, both on average and in the minority group. Data-centric approaches such as UV-DRO (Srivastava et al., 2020) and LISA (Yao et al., 2022) assume full information about the type of the spurious correlations on all instances and augment the training data with additional samples to eliminate the spurious correlation. GroupDRO (Sagawa et al., 2020a) instead uses the existing group labels directly to minimize the loss for the worst-group performance. Requiring only a small subset of group labeled training data, Barack (Sohoni et al., 2022) predicts other instances’ groups and DFR (Kirichenko et al., 2023; Izmailov et al., 2022) retrains part of the model with a small, spurious-free dataset. Group annotation or human-in-the-loop annotations is, however, costly and difficult to achieve in practice. In particular complex spurious attributes, such as as fluency bias (Wang et al., 2023) or texture and shape bias (Geirhos et al., 2018) remain challenging, as they are difficult for humans to detect.
@@
Assuming no group annotations, Sohoni et al. (2020); Seo et al. (2022) and Creager et al. (2021) focus on automatically inferring appropriate groups before training a robust model (e.g., GroupDRO) with pseudo-group labels. Several methods follow a two-stage training approach, where the first stage is commonly used to identify hard cases from ERM training. These hard cases are instances that the ERM model fails to classify, presumably because ERM relies on spurious correlations, with the hard-to-classify cases not exhibiting the spurious correlation, i.e., bias-conflicting samples. Not utilizing the misclassified cases, SPARE (Yang et al., 2024) infers groups labels by clustering model outputs in early epochs, then applies importance sampling to upsample smaller clusters (minority groups) and downsample larger clusters (majority groups) during the remaining epochs. Approaches differ in the second stage: Liu et al. (2021); Nam et al. (2020) upweight hard cases in an additional training run, while Yaghoobzadeh et al. (2021) exclusively fine-tune hard cases. CnC (Zhang et al., 2022) and DCWP (Park et al., 2023) utilize the set of hard cases for contrastive learning to define contrastive batches and retrain the whole network (CnC) or to extract a subnetwork (DCWP) with contrastive loss. Different from the two-stage approaches, DeDiER (Tiwari et al., 2024) uses knowledge distillation to find overconfident predictions in early network layers $d$, and then uses this signal to weight the distillation loss on a per-instance level. SELF (LaBonte et al., 2024) finds a small set of hard cases by measuring the degree of disagreement in predictions between an ERM and an early-stop ERM model, and then applies last-layer fine-tuning with this set. The effectiveness of methods that rely on ERM misclassifications for hard cases depends on both the quantity and the quality of the misclassifications. In cases where ERM models almost perfectly fit the training data (training accuracy close to $100\%$), there are not enough hard cases (i.e., misclassifications) and these methods fail. To obtain a sufficiently large set of hard cases, methods either use a carefully controlled early stopping criterion (Liu et al., 2021) or classify based on earlier layers in the networks, assuming that these layers still contain information about misclassified samples, likely originate from learned spurious correlations (DeDiER (Tiwari et al., 2024)). Therefore, methods that rely on misclassification cannot directly mitigate spurious correlations of a trained ERM model unless trainers have access to the entire training process of that model.
@@ 3 Problem Setting and Example
In this section, we describe spurious correlations in machine learning more formally (Sec.3.1), and show the core idea of our method PruSC for mitigating spurious correlations in a simple example (Sec.3.2).
@@ 3.1 Problem Setting
Suppose a data set $\mathcal{D}$ consists of data where each sample has some attributes of the input $x$ that are correlated with the target labels $y\in\mathcal{Y}$. With prior knowledge of spurious attributes $a\in\mathcal{A}$, we can partition the data set $\mathcal{D}$ into groups $g\in\mathcal{G}$ according to $\mathcal{G}=\mathcal{A}\times\mathcal{Y}$, i.e., each group $g$ is defined by the combination of the label and the corresponding spurious attribute (Sagawa et al., 2020b).
A model trained with Empirical Risk Minimization (ERM) on the class labels may be able to separate the classes well but rely on spurious attributes to do so, resulting in a mismatch between the intended and the model-learned solution (Geirhos et al., 2020). Spurious attributes typically appear in the majority of the dataset. Thus, by using these attributes, an ERM model can minimize the training loss for the expected average population, but still show high errors on the minority group that lacks these spurious attributes. Furthermore, in cases where the spurious attributes are easy to learn (e.g., background instead of the object), ERM prioritizes learning those attributes over more complex ones and then loses the ability to generalize. Our goal is to mitigate the learning of spurious correlations by addressing both statistical (highly unbalanced group distributions) and geometric (the tendency to learn simpler features) issues.
@@ 3.2 Motivating Example
To illustrate the problem and motivate our approach, we consider a simple 2-D classification task using the two moons dataset (Pezeshki et al., 2021), with one attribute that is strongly correlated with target labels, causing neural networks to under-learn other invariant attributes. The two moons dataset $\mathcal{D}$ consists of input pairs $(x,y)$, where $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ and $y\in\{0,1\}$ (cf. Fig. 1). The dataset contains a strong relation between $x_{1}$ and the corresponding label $y$. We train three different models on this dataset, showing the effect of i) smaller networks and ii) our contrastive loss formulation (Sec. 5.2.1). A simple feed-forward network (5 ReLu layers, 500 hidden units per layer, 100 epochs, cross-entropy loss) can separate the two classes very well (Fig. 1, left). The decision boundary is nearly linear because the cross-entropy loss optimizes for class separation and for the model it is sufficient to rely on the correlation between the x-axis values of $x$ and $y$ to minimize the loss. We train the same network architecture with random masks (50% of weights in each layer) and use the same fraction of random weights for testing, i.e., we randomly prune the network to 50% of its weights. We observe curved decision boundaries (Fig. 1, center), suggesting that a smaller network within a dense network is more robust to invariant features. Combining masked training and our contrastive loss (Sec. 5.2.1, we obtain a curved decision boundary with a large margin (Fig. 1, right), suggesting that the model uses information from both the x- and y-coordinates.
@@
Figure 1: Simple network on the two moons dataset. Decision boundaries (black lines) depend on model capacity and loss. Left: Standard ERM with cross-entropy loss with a nearly linear decision boundary shows a strong dependence on the x-coordinate (nearly linear vertical decision boundary). Center: Pruned network (masking) using only 50% of the weights at test time shows less dependence on the x-coordinate. Right: Masking, and using our contrastive loss (cf. Sec. 5.2.1) results in non-linear decision boundaries with large margins.
@@
The above example strengthens two hypotheses for mitigating spurious correlations. First, network capacity seems to have an influence. Both, the dense network and the smaller subnetwork with randomly pruned connections perform well on the dataset, but the pruned network has better decision boundaries, i.e., depends less on the spurious feature. However, this does not suggest to train a sparse architecture from scratch, because small architectures are difficult to train, whereas larger models have a higher ability to learn information and improve performance beyond the bias-variance regime (Nakkiran et al., 2021; Frankle & Carbin, 2018). Instead, it suggests to first train a large network with sufficient capacity to encode all relevant information, and then extract a subnetwork after training that does not rely on spurious correlations. The post-training approach is also consistent with our realistic scenario that the existence of spurious correlations becomes apparent only after training (and possibly later in the lifetime of a model). Thus, extracting a subnetwork without spurious correlations is less expensive than retraining the whole model from scratch. This observation is our main motivation. Second, our contrastive learning scheme improves decision boundaries by learning more information from the training data than just the simplicity bias features. Our batches of contrastive instances serve as a guide for the models to know which feature relations to focus less on.